40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3578     Entry time: Wed Sep 15 16:12:35 2010
Author: koji, steve 
Type: Update 
Category: MOPA 
Subject: MOPA Controller is taken out of the PSL rack 

We removed the Lightwave MOPA Controller, PA#102, NPRO206 power supply to make room for the IOO chassy at 1X1 (south) rack.

The umbilical cord was a real pain to take out. It is shading its plastic cover. The unused Minco was disconnected and removed.

The ref. cavity ion pump controller- power supply was temporarily taken out also.

Attachment 1: P1060843.JPG  2.552 MB  Uploaded Wed Sep 15 17:13:40 2010  | Hide | Hide all
P1060843.JPG
ELOG V3.1.3-